# Learning to Solve Integer Linear Programs with Davis-Yin Splitting

### v3

@@ Abstract
In many applications, a combinatorial problem must be repeatedly solved with similar, but distinct parameters. Yet, the parameters $w$ are not directly observed; only contextual data $d$ that correlates with $w$ is available. It is tempting to use a neural network to predict $w$ given $d$. However, training such a model requires reconciling the discrete nature of combinatorial optimization with the gradient-based frameworks used to train neural networks. When the problem in question is an Integer Linear Program (ILP), one approach to overcome this training issue is to consider a continuous relaxation of the combinatorial problem. While existing methods utilizing this approach have shown to be highly effective on small problems, they do not always scale well to large problems. In this work, we draw on ideas from modern convex optimization to design a network and training scheme which scales effortlessly to problems with thousands of variables. Our experiments verify the computational advantage our proposed method enjoys on two representative problems, namely the shortest path problem and the knapsack problem.
@@ 1 Introduction
Many high-stakes decision problems in healthcare Zhong & Tang (2021), logistics and scheduling Sbihi & Eglese (2010); Kacem et al. (2021), and transportation Wang & Tang (2021) can be viewed as a two step process. In the first step, one gathers data about the situation at hand. This data is used to assign a value (or cost) to outcomes arising from each possible action. The second step is to select the action yielding maximum value (alternatively, lowest cost). Mathematically, this can be framed as an optimization problem with a data-dependent cost function:
@@
In this work, we focus on the case where $\mathcal{X}\subset\mathbb{R}^{n}$ is a finite constraint set and $f(x;d)=w(d)^{\top}x$ is a linear function. This class of problems is quite rich, containing the shortest path, traveling salesperson, and sequence alignment problems, to name a few. Given $w(d)$, solving (1) may be straightforward (e.g. shortest path) or NP-hard (e.g. traveling salesperson problem Karp (1972)). However, our present interest is settings where the dependence of $w(d)$ on $d$ is unknown. In such settings, it is intuitive to learn a mapping $w_{\Theta}(d)\approx w(d)$ and then solve
@@
At first glance, it may appear natural to tune the weights $\Theta$ so as to minimize the difference between $w_{\Theta}(d)$ and $w(d)$. However, this is only possible if $w(d)$ is available at training time. Even if this approach is feasible, it may not be advisable. If $w_{\Theta}(d)$ is a near-perfect predictor of $w(d)$ we can expect $x_{\Theta}(d)\approx x^{\star}(d)$. However, if there are even small differences between $w_{\Theta}(d)$ and $w(d)$ this can manifest in wildly different solutions Bengio (1997); Wilder et al. (2019). Thus, we focus on methods where $\Theta$ is tuned by minimizing the discrepancy between $x_{\Theta}(d)$ and $x^{\star}(d)$.
The key obstacle in using gradient-based algorithms to tune $\Theta$ in such approaches is “differentiating through” the solution $x_{\Theta}(d)$ of (2) to obtain a gradient with which to update $\Theta$. Specifically, the combinatorial nature of $\mathcal{X}$ may cause the solution ${x}_{\Theta}(d)$ to remain unchanged for many small perturbations to $\Theta$; yet, for some perturbations ${x}_{\Theta}(d)$ may “jump” to a different point in $\mathcal{X}$. Hence, the gradient $\mbox{d}{x}_{\Theta}\big{/}\mbox{d}w_{\Theta}$ is always either zero or undefined. To compute an informative gradient, we follow Wilder et al. (2019) by relaxing (2) to a quadratic program over the convex hull of $\mathcal{X}$ with an added regularizer (see (8)).
@@
Figure 1: The shortest path prediction problem Pogančić et al. (2019). The goal is to find the shortest path (from top-left to bottom-right) through a randomly generated terrain map from the Warcraft II tileset Guyomarch . The contextual data $d$, shown in (a), is an image sub-divided into 8-by-8 squares, each representing a vertex in a 12-by-12 grid graph. The cost of traversing each square, i.e. $w(d)$, is shown in (b), with darker shading representing lower cost. The true shortest path is shown in (c).
@@
Drawing upon recent advances in convex optimization Ryu & Yin (2022) and implicit neural networks Fung et al. (2022); McKenzie et al. (2021), we propose a method designed specifically for “differentiating through” large-scale integer linear programs. Our approach is fast, easy to implement using our provided code, and, unlike several prior works (e.g. see Pogančić et al. (2019); Berthet et al. (2020)), trains completely on GPU. Numerical examples herein demonstrate our approach, run using only standard computing resources, easily scales to problems with tens of thousands of variables. Theoretically, we verify our approach computes an informative gradient via a refined analysis of Jacobian-free Backpropagation (JFB) Fung et al. (2022). In summary, we do the following.
@@ 2 Preliminaries
We focus on optimization problems of the form (1) where $f(x;d)=w(d)^{\top}x$ and $\mathcal{X}$ is the integer or binary points of a polytope, which we may assume to be expressed in standard form Ziegler (2012):
@@
In other words, (1) is an Integer Linear Program (ILP). Similar to works by Elmachtoub & Grigas (2022); Wilder et al. (2019); Mandi & Guns (2020) and others we replace the model (2) with its continuous relaxation, and redefine
@@
We assume access to training data in the tuple form $(d,x^{\star}(d))$. We focus on the $\ell_{2}$ loss:
@@
and $\mathcal{D}$ is the distribution of contextual data, although more sophisticated losses can be used with only superficial changes to our results. In principle we select the optimal weights by solving $\operatorname*{arg\,min}_{\Theta}\mathcal{L}(\Theta)$. In practice, the population risk is inaccessible, and so we minimize empirical risk instead Vapnik (1999):
@@
As discussed in Section 1, ${x}^{\star}_{\Theta}$ is piecewise constant as a function of $w_{\Theta}$, and this remains true for the LP relaxation (4). Consequently, for all $w_{\Theta}$, either $\partial{x}_{\Theta}/\partial\Theta=0$ or $\partial{x}_{\Theta}/\partial\Theta$ is undefined—neither case yields an informative gradient. To remedy this, Wilder et al. (2019); Mandi & Guns (2020) propose adding a small amount of regularization to the objective function in (4) to make it strongly convex. This ensures ${x}_{\Theta}$ is a continuously differentiable function of $w_{\Theta}$. We add a small quadratic regularizer, modulated by $\gamma\geq 0$, to obtain the objective
@@
During training, we aim to solve (8) and efficiently compute the derivative $\partial{x}_{\Theta}/\partial\Theta$.
@@ 3 Related Works
We highlight recent works combining optimization—particularly mixed integer programs—and deep learning.
@@
Deep Equilibrium Models
Bai et al. (2019); El Ghaoui et al. (2021) propose the notion of deep equilibrium model (DEQ), also known as an implicit neural network. A DEQ is a neural network for which forward propagation consists of (approximately) computing a fixed point of a parametrized operator. We note that (8) may be reformulated as a fixed point problem,
where $P_{\mathcal{C}}$ is the orthogonal projection [^1] For a set $\mathcal{A}\subseteq{\mathbb{R}}^{n}$, $P_{\mathcal{A}}(x)\triangleq\operatorname*{arg\,min}_{z\in\mathcal{A}}\|z-x\|$. onto $\mathcal{C}$. Thus, DEQ techniques may be applied to constrained optimization layers Chen et al. (2021); Blondel et al. (2022); McKenzie et al. (2021). However, the cost of computing $P_{\mathcal{C}}$ can be prohibitive, see the discussion in Section 4.
Differentiable Combinatorial Optimizers
We now narrow our focus to optimization layers for (integer) linear programs. Wilder et al. (2019) and Elmachtoub & Grigas (2022) are two of the earliest works applying deep learning techniques to data-driven LPs, and delineate two fundamentally different approaches. Specifically, Wilder et al. (2019) proposes replacing the LP with a continuous and strongly convex relaxation, as described in Section 2. This approach is extended to ILPs in Ferber et al. (2020), and to non-quadratic regularizers in Mandi & Guns (2020). On the other hand, Elmachtoub & Grigas (2022) avoid modifying the underlying optimization problem and instead propose a new loss function; dubbed the Smart Predict-then-Optimize (SPO) loss, to be used instead of the $\ell_{2}$ loss. We emphasize that the SPO loss requires access to the true cost vectors $w(d)$, a setting which we do not consider.
Several works define a continuously differentiable proxy for the solution to the unregularized LP (4), which we rewrite here as [^2] For sake of notational simplicity, the dependence on $d$ is implicit.
@@
which is somewhat analogous to Nesterov-Spokoiny smoothing Nesterov & Spokoiny (2017) in zeroth-order optimization. Pogančić et al. (2019) define a piecewise-affine interpolant to $\ell(x^{\star}(w))$, where $\ell$ is a suitable loss function.
We note a bifurcation in the literature; Wilder et al. (2019); Ferber et al. (2020); Elmachtoub & Grigas (2022) assume access to training data of the form $(d,w(d))$, whereas Pogančić et al. (2019); Berthet et al. (2020); Sahoo et al. (2022) use training data of the form $(d,x^{\star}(d))$. Our focus is on the latter setting, and we refer the reader to Kotary et al. (2021) for a recent survey of this area.
@@
Our work can be situated within the broader learning-to-optimize (L2O) framework Chen et al. (2022); Li & Malik (2017); Chen et al. (2018); Liu et al. (2023), where an optimization algorithm is learned and its outputs are used to perform inferences. However, traditional L2O methods use a fixed number of layers (i.e. unrolled algorithms). Our approach attempts to differentiate through the solution $x^{\star}_{\Theta}$ and is therefore most closely aligned with works that employ implicit networks within the L2O framework Amos & Kolter (2017); Heaton & Fung (2023); McKenzie et al. (2021); Gilton et al. (2021); Liu et al. (2022). We highlight that, unlike many L20 works, our goal is not to learn a faster optimization method for fully specified problems. Rather, we seek to solve partially specified problems given contextual data by combining learning and optimization techniques.
@@
where $\hat{\lambda}$ and $\hat{\nu}\geq 0$ are Lagrange multipliers associated to the optimal solution ${x}_{\Theta}$ Bertsekas (1997) and $D(\hat{\nu})$ is a matrix with $\hat{\nu}$ along its diagonal. Differentiating these equations with respect to $\Theta$ and rearranging yields
@@
The matrix and right hand side vector in (12) are computable, thus enabling one to solve for $\frac{d{x}_{\Theta}}{d\Theta}$ (as well as $\frac{d\hat{\lambda}}{d\Theta}$ and $\frac{d\hat{\nu}}{d\Theta}$). The computational bottleneck in this approach is computing the Lagrange multipliers at optimality—i.e. $\hat{\lambda}$ and $\hat{\nu}$—in addition to ${x}_{\Theta}$. If $x\in{\mathbb{R}}^{n}$ and $A\in\mathbb{R}^{m\times n}$ this can be done with a primal-dual interior point method at a cost of $\mathcal{O}\left(\max\{n,m\}^{3}\right)$ Amos & Kolter (2017).
Using the stochastic proxy (11), Berthet et al. (2020) derive a formula for $dx^{\star}_{\varepsilon}/dw$ which is also an expectation, and hence can be efficiently approximated using Monte Carlo methods. Pogančić et al. (2019) show that the gradients of their interpolant are strikingly easy to compute, requiring just one additional solve of (10) with perturbed cost $w^{\prime}$. This approach is extended by Sahoo et al. (2022) which proposes to avoid computing $dx^{\star}_{\varepsilon}/dw$ entirely, replacing it with the negative identity matrix. This is similar in spirit to our use of Jacobian-free Backpropagation (see Section 4).

### v4

# Differentiating Through Integer Linear Programs with Quadratic Regularization and Davis-Yin Splitting

@@ Abstract
In many applications, a combinatorial problem must be repeatedly solved with similar, but distinct parameters. Yet, the parameters $w$ are not directly observed; only contextual data $d$ that correlates with $w$ is available. It is tempting to use a neural network to predict $w$ given $d$. However, training such a model requires reconciling the discrete nature of combinatorial optimization with the gradient-based frameworks used to train neural networks. We study the case where the problem in question is an Integer Linear Program (ILP). We propose applying a three-operator splitting technique, also known as Davis-Yin splitting (DYS), to the quadratically regularized continuous relaxation of the ILP. We prove that the resulting scheme is compatible with the recently introduced Jacobian-free backpropagation (JFB). Our experiments on two representative ILPs: the shortest path problem and the knapsack problem, demonstrate that this combination—DYS on the forward pass, JFB on the backward pass—yields a scheme which scales more effectively to high-dimensional problems than existing schemes. All code associated with this paper is available at https://github.com/mines-opt-ml/fpo-dys.
@@ 1 Introduction
Many high-stakes decision problems in healthcare (Zhong & Tang, 2021), logistics and scheduling (Sbihi & Eglese, 2010; Kacem et al., 2021), and transportation (Wang & Tang, 2021) can be viewed as a two step process. In the first step, one gathers data about the situation at hand. This data is used to assign a value (or cost) to outcomes arising from each possible action. The second step is to select the action yielding maximum value (alternatively, lowest cost). Mathematically, this can be framed as an optimization problem with a data-dependent cost function:
@@
In this work, we focus on the case where $\mathcal{X}\subset\mathbb{R}^{n}$ is a finite constraint set and $f(x;d)=w(d)^{\top}x$ is a linear function. This class of problems is quite rich, containing the shortest path, traveling salesperson, and sequence alignment problems, to name a few. Given $w(d)$, solving equation 1 may be straightforward (e.g. shortest path) or NP-hard (e.g. traveling salesperson problem (Karp, 1972)). However, our present interest is settings where the dependence of $w(d)$ on $d$ is unknown. In such settings, it is intuitive to learn a mapping $w_{\Theta}(d)\approx w(d)$ and then solve
@@
At first glance, it may appear natural to tune the weights $\Theta$ so as to minimize the difference between $w_{\Theta}(d)$ and $w(d)$. However, this is only possible if $w(d)$ is available at training time. Even if this approach is feasible, it may not be advisable. If $w_{\Theta}(d)$ is a near-perfect predictor of $w(d)$ we can expect $x_{\Theta}(d)\approx x^{\star}(d)$. However, if there are even small differences between $w_{\Theta}(d)$ and $w(d)$ this can manifest in wildly different solutions (Bengio, 1997; Wilder et al., 2019). Thus, we focus on methods where $\Theta$ is tuned by minimizing the discrepancy between $x_{\Theta}(d)$ and $x^{\star}(d)$. Such methods are instances of the decision-focused learning paradigm Mandi et al. (2023), as the criterion we are optimizing for is the quality of $x_{\Theta}(d)$ (the “decision”) not the fidelity of $w_{\Theta}(d)$ (the “prediction”).
The key obstacle in using gradient-based algorithms to tune $\Theta$ in such approaches is “differentiating through” the solution $x_{\Theta}(d)$ of equation 2 to obtain a gradient with which to update $\Theta$. Specifically, the combinatorial nature of $\mathcal{X}$ may cause the solution ${x}_{\Theta}(d)$ to remain unchanged for many small perturbations to $\Theta$; yet, for some perturbations ${x}_{\Theta}(d)$ may “jump” to a different point in $\mathcal{X}$. Hence, the gradient $\mbox{d}{x}_{\Theta}\big{/}\mbox{d}w_{\Theta}$ is always either zero or undefined. To compute an informative gradient, we follow Wilder et al. (2019) by relaxing equation 2 to a quadratic program over the convex hull of $\mathcal{X}$ with an added regularizer (see equation 8).
@@
Figure 1: The shortest path prediction problem (Pogančić et al., 2019). The goal is to find the shortest path (from top-left to bottom-right) through a randomly generated terrain map from the Warcraft II tileset (Guyomarch, ). The contextual data $d$, shown in (a), is an image sub-divided into 8-by-8 squares, each representing a vertex in a 12-by-12 grid graph. The cost of traversing each square, i.e. $w(d)$, is shown in (b), with darker shading representing lower cost. The true shortest path is shown in (c).
@@
Drawing upon recent advances in convex optimization (Ryu & Yin, 2022) and implicit neural networks (Fung et al., 2022; McKenzie et al., 2021), we propose a method designed specifically for “differentiating through” large-scale integer linear programs. Our approach is fast, easy to implement using our provided code, and, unlike several prior works (e.g. see Pogančić et al. (2019); Berthet et al. (2020)), trains completely on GPU. Numerical examples herein demonstrate our approach, run using only standard computing resources, easily scales to problems with tens of thousands of variables. Theoretically, we verify our approach computes an informative gradient via a refined analysis of Jacobian-free Backpropagation (JFB) (Fung et al., 2022). In summary, we do the following.
@@ 2 Preliminaries
We focus on optimization problems of the form equation 1 where $f(x;d)=w(d)^{\top}x$ and $\mathcal{X}$ is the integer or binary points of a polytope, which we may assume to be expressed in standard form (Ziegler, 2012):
@@
In other words, equation 1 is an Integer Linear Program (ILP). Similar to works by Elmachtoub & Grigas (2022); Wilder et al. (2019); Mandi & Guns (2020) and others we replace the model equation 2 with its continuous relaxation, and redefine
@@
We assume access to training data in the tuple form $(d,x^{\star}(d))$. Such data may be obtained by observing experienced agents solve equation 1, for example taxi drivers in a shortest path problem Piorkowski et al. (2009). We focus on the $\ell_{2}$ loss:
@@
and $\mathcal{D}$ is the distribution of contextual data, although more sophisticated losses can be used with only superficial changes to our results. In principle we select the optimal weights by solving $\operatorname*{arg\,min}_{\Theta}\mathcal{L}(\Theta)$. In practice, the population risk is inaccessible, and so we minimize empirical risk instead (Vapnik, 1999):
@@
As discussed in Section 1, ${x}^{\star}_{\Theta}$ is piecewise constant as a function of $w_{\Theta}$, and this remains true for the LP relaxation equation 4. Consequently, for all $w_{\Theta}$, either $\partial{x}_{\Theta}/\partial\Theta=0$ or $\partial{x}_{\Theta}/\partial\Theta$ is undefined—neither case yields an informative gradient. To remedy this, Wilder et al. (2019); Mandi & Guns (2020) propose adding a small amount of regularization to the objective function in equation 4 to make it strongly convex. This ensures ${x}_{\Theta}$ is a continuously differentiable function of $w_{\Theta}$. We add a small quadratic regularizer, modulated by $\gamma\geq 0$, to obtain the objective
@@
During training, we aim to solve equation 8 and efficiently compute the derivative $\partial{x}_{\Theta}/\partial\Theta$.
Pitfalls of Relaxation
Finally, we offer a note of caution when using the quadratically regularized problem equation 8 as a proxy for the original integer linear program equation 1. Temporarily defining
$\hat{x}_{\Theta}(d)\triangleq\operatorname*{arg\,min}_{x\in\mathcal{X}}w_{\Theta}(d)^{\top}x\quad\tilde{x}_{\Theta}(d)\triangleq\operatorname*{arg\,min}_{x\in\mathcal{C}}w_{\Theta}(d)^{\top}x\quad{x}_{\Theta}(d)\triangleq\operatorname*{arg\,min}_{x\in\mathcal{C}}f_{\Theta}(x;\gamma,d),$
we observe that two sources of error have arisen; one from replacing $\mathcal{X}$ with $\mathcal{C}$ (i.e. the discrepancy between $\hat{x}_{\Theta}$ and $\tilde{x}_{\Theta}$), and one from replacing the linear objective function with its quadratically regularized counterpart (i.e. the discrepancy between $\tilde{x}_{\Theta}$ and $x_{\Theta}$). When $A$ is totally unimodular, $\tilde{x}_{\Theta}$ is guaranteed to be integral, in which case $\hat{x}_{\Theta}=\tilde{x}_{\Theta}$. Moreover, Wilder et al. (2019, Theorem 2) shows that
$0\leq w_{\Theta}(d)^{\top}\left(x_{\Theta}(d)-\tilde{x}_{\Theta}(d)\right)\leq\gamma\max_{x,y\in\mathcal{C}}\|x-y\|^{2}$
whence, as in the proof Berthet et al. (2020, Proposition 2.2), $\lim_{\gamma\to 0^{+}}x_{\Theta}(d)=\tilde{x}_{\Theta}(d)$ (at least when $\mathcal{C}$ is compact). Thus, for totally unimodular ILPs (e.g. the shortest path problem), it is reasonable to expect that $x_{\Theta}(d)$ obtained from equation 8, for well-tuned $\Theta$ and small $\gamma$, are of acceptable quality. For ILPs which are not totally unimodular no such theoretical guarantees exist. However, we observe that this relaxation works reasonably well, at least for the knapsack problem.
@@ 3 Related Works
We highlight recent works in decision-focused learning—particularly for mixed integer programs—and related areas.
@@
Decision-focused learning for LPs
When an optimization layer is fitted to a data-dependent optimization problems of the form equation 1, with the goal of maximizing the quality of the predicted solution, this is usually referred to as decision-focused learning. We narrow our focus to the case where the objective function in equation 1 is linear.Wilder et al. (2019) and Elmachtoub & Grigas (2022) are two of the earliest works applying deep learning techniques to data-driven LPs, and delineate two fundamentally different approaches. Specifically, Wilder et al. (2019) proposes replacing the LP with a continuous and strongly convex relaxation, as described in Section 2. This approach is extended to ILPs in Ferber et al. (2020), and to non-quadratic regularizers in Mandi & Guns (2020). On the other hand, Elmachtoub & Grigas (2022) avoid modifying the underlying optimization problem and instead propose a new loss function; dubbed the Smart Predict-then-Optimize (SPO) loss, to be used instead of the $\ell_{2}$ loss. We emphasize that the SPO loss requires access to the true cost vectors $w(d)$, a setting which we do not consider.
Several works define a continuously differentiable proxy for the solution to the unregularized LP equation 4, which we rewrite here as [^1] For sake of notational simplicity, the dependence on $d$ is implicit.
@@
which is somewhat analogous to Nesterov-Spokoiny smoothing (Nesterov & Spokoiny, 2017) in zeroth-order optimization. For some draws of $Z$ the additively perturbed cost vector $w+\varepsilon z$ may have negative entries, which may cause a solver (e.g. Dijkstra’s algorithm) applied to the associated LP to fail. To avoid this, Dalle et al. (2022) suggest using a multiplicative perturbation of $w$ instead. Pogančić et al. (2019) define a piecewise-affine interpolant to $\ell(x^{\star}(w))$, where $\ell$ is a suitable loss function.
We note a bifurcation in the literature; Wilder et al. (2019); Ferber et al. (2020); Elmachtoub & Grigas (2022) assume access to training data of the form $(d,w(d))$, whereas Pogančić et al. (2019); Berthet et al. (2020); Sahoo et al. (2022) use training data of the form $(d,x^{\star}(d))$. The former setting is frequently referred to as the predict-then-optimize problem. Our focus is on the latter setting. We refer the reader to Kotary et al. (2021); Mandi et al. (2023) for recent surveys of this area.
Deep Equilibrium Models
Bai et al. (2019); El Ghaoui et al. (2021) propose the notion of deep equilibrium model (DEQ), also known as an implicit neural network. A DEQ is a neural network for which forward propagation consists of (approximately) computing a fixed point of a parametrized operator. We note that equation 8 may be reformulated as a fixed point problem,
where $P_{\mathcal{C}}$ is the orthogonal projection [^2] For a set $\mathcal{A}\subseteq\mathbb{R}^{n}$, $P_{\mathcal{A}}(x)\triangleq\operatorname*{arg\,min}_{z\in\mathcal{A}}\|z-x\|$. onto $\mathcal{C}$. Thus, DEQ techniques may be applied to constrained optimization layers (Chen et al., 2021; Blondel et al., 2022; McKenzie et al., 2021). However, the cost of computing $P_{\mathcal{C}}$ can be prohibitive, see the discussion in Section 4.
@@
Our work is related to the learning-to-optimize (L2O) framework (Chen et al., 2022; Li & Malik, 2017; Chen et al., 2018; Liu et al., 2023), where an optimization algorithm is learned and its outputs are used to perform inferences. However, traditional L2O methods use a fixed number of layers (i.e. unrolled algorithms). Our approach attempts to differentiate through the solution $x^{\star}_{\Theta}$ and is therefore most closely aligned with works that employ implicit networks within the L2O framework (Amos & Kolter, 2017; Heaton & Fung, 2023; McKenzie et al., 2021; Gilton et al., 2021; Liu et al., 2022). We highlight that, unlike many L2O works, our goal is not to learn a faster optimization method for fully specified problems. Rather, we seek to solve partially specified problems given contextual data by combining learning and optimization techniques.
@@
where $\hat{\lambda}$ and $\hat{\nu}\geq 0$ are Lagrange multipliers associated to the optimal solution ${x}_{\Theta}$ (Bertsekas, 1997) and $D(\hat{\nu})$ is a matrix with $\hat{\nu}$ along its diagonal. Differentiating these equations with respect to $\Theta$ and rearranging yields
@@
The matrix and right hand side vector in equation 12 are computable, thus enabling one to solve for $\frac{d{x}_{\Theta}}{d\Theta}$ (as well as $\frac{d\hat{\lambda}}{d\Theta}$ and $\frac{d\hat{\nu}}{d\Theta}$). We emphasize that both primal (i.e., $x_{\Theta}$) and dual (i.e., $\hat{\lambda}$ and $\hat{\nu}$) variables at optimality are required. This constrains the choice of optimization schemes to those that track primal and dual variables, and prohibits the use of fast, primal-only schemes. Following (Amos & Kolter, 2017), most works employing this approach use a primal-dual interior point method, which computes acceptable approximations to $x_{\Theta}$, $\hat{\lambda}$, and $\hat{\nu}$ at a cost of $\mathcal{O}\left(\max\{n,m\}^{3}\right)$, assuming $x\in\mathbb{R}^{n}$ and $A\in\mathbb{R}^{m\times n}$.
Using the stochastic proxy equation 10, Berthet et al. (2020) derive a formula for $dx^{\star}_{\varepsilon}/dw$ which is also an expectation, and hence can be efficiently approximated using Monte Carlo methods. Pogančić et al. (2019) show that the gradients of their interpolant are strikingly easy to compute, requiring just one additional solve of equation 9 with perturbed cost $w^{\prime}$. This approach is extended by Sahoo et al. (2022) which proposes to avoid computing $dx^{\star}_{\varepsilon}/dw$ entirely, replacing it with the negative identity matrix. This is similar in spirit to our use of Jacobian-free Backpropagation (see Section 4).